OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1 (1, 5) (-5, -1) (5, 1) (-5, 1)

OpenStudy (abb0t):

Use algebra to rearrange one of the equations. I suggest rearranging equation number 2. and substituting for "x". Since all you have to do is add one term.

OpenStudy (anonymous):

Find the value of x from second equation first.. can you do that??

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

x - 4y = 1 Here, add 4y both the sides and tell me what will you get??

OpenStudy (anonymous):

x8=5?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

You cannot do like this: 1 + 4y = 5 (wrong).. 1 and 4y are unlike terms you cannot add or subtract them..

OpenStudy (anonymous):

so I don't know what to do? @waterineyes

OpenStudy (anonymous):

See : \[x - 4y = 1\] Add 4y both the sides: \[x - 4y + 4y = 1 + 4y\] \[x - \cancel{4y} + \cancel{4y} = 1 + 4y\] \[x = 4y + 1\] Now put this value of x in first equation to get y..

OpenStudy (anonymous):

so is it B?

OpenStudy (anonymous):

Solved or making guess??

OpenStudy (anonymous):

Try more...

OpenStudy (anonymous):

guess

OpenStudy (anonymous):

@waterineyes come on! I'm begging you here

OpenStudy (abb0t):

You gotta put in the work if you want to learn and master the material.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin4.htm

OpenStudy (anonymous):

\[2x - 9y = 1\]\[x - 4y = 1\] \[x=4y+1\]\[2(4y+1)-9y=1\]\[8y+2-9y=1\]\[-y=-1\]\[y=1\] \[2x-9(1)=1\]\[2x=10\]\[x=5\] \[\large (5,1)\] grow the heck up and actually learn how to do it instead of expecting others to do it for you

OpenStudy (anonymous):

I'm sorry? I have trouble in math and I have to take special classes to help me. A person can't learn everything buddy.

OpenStudy (anonymous):

you can if you want to, you just choose not to

OpenStudy (anonymous):

I just have trouble in math. I try to learn but I always seem to fail for some reason when it comes to math.

OpenStudy (anonymous):

if i, who has only a few more years to live, can change my attitude even though i have very special needs and force myself to learn this stuff and keep up a 5.0 GPA even though i wont even live long enough to be able to use that in life, you can at least try, I'm very sure..

OpenStudy (anonymous):

that sucks... but not everyone has a mind set like you do. Sure you can try, but you're not always going to succeed.

OpenStudy (anonymous):

again, not with that attitude..

OpenStudy (anonymous):

@yummydum what has happened to you??

OpenStudy (anonymous):

i dont share :)

OpenStudy (anonymous):

C'mon tell me buddy..

OpenStudy (anonymous):

God wants me back early, that's all i tell :)

OpenStudy (anonymous):

How old are you?

OpenStudy (anonymous):

How old are you??

OpenStudy (anonymous):

wow. we asked the same question... lol

OpenStudy (anonymous):

What a coincidence..

OpenStudy (anonymous):

Ha ha ha ha...

OpenStudy (anonymous):

what! that sucks! sorry to hear that you aren't going to live that long

OpenStudy (anonymous):

Tell me in inbox, I will not share it with any other...

OpenStudy (anonymous):

thats a wonderful way to put it christian, wow lol

OpenStudy (anonymous):

sorry.. I don't know how else to say it:/

OpenStudy (anonymous):

@waterineyes its not a condition its just that i was born to only live so long :) God planned an early death for me, that is all, nothing to worry about :)

OpenStudy (anonymous):

There must be any problem.. And please Don't blame God..

OpenStudy (anonymous):

I will not force you, and I pray that you Live Long..

OpenStudy (anonymous):

He's not blamming God, he's just stating that God wants him back.

OpenStudy (anonymous):

He or she is mentioning God every time.. That sounds like Blaming..

OpenStudy (anonymous):

oh lol sorry(: she*

OpenStudy (anonymous):

blaming would be if i said "its God's fault" i said "this is what God planned for me :) and I accept it"

OpenStudy (anonymous):

I don't think she's blamming God

OpenStudy (anonymous):

Okay Okay, Sorry...

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

well you guys have a good one :)

OpenStudy (anonymous):

you too!